Suppose the government applies a specific tax to a good where the demand​ elasticity, epsilon​, is negative 0.9​, and the supply​ elasticity, eta​, is 0.9. If a specific​ tax, tau​, of ​$1.75 was placed on the​ good, to the nearest​ cent, what is the price increase that consumers would​ pay? To the nearest​ cent, what is the price decrease that producers would​ pay? ​(round your answer to two decimal places​). What is the tax incidence on​ consumers? ​(round your answer to two decimal places​).
Answer #1

Burden of tax on consumers = eta/(eta + epsilon) = 0.9/(0.9+0.9) = 0.9/1.8 = 1/2
So, price increase that consumers would​ pay = Burden of tax on consumers*amount of tax = (1/2)*1.75 = 0.88 cents

Burden of tax on producers = epsilon/(eta+epsilon) = 0.9/(0.9+0.9) = 0.9/1.8 = 1/2
So, price decrease that producers would​ pay = Burden of tax on producers*amount of tax = (1/2)*1.75 = 0.88 cents

Tax incidence on consumers = 1/2 = 0.50 of the tax

(Note: Absolute value of epsilon is taken.)